OpenStudy (anonymous):

solving an equation using zero product property (x-2) (x+7)=0 I think it's 2,-7 but i am not certain.

OpenStudy (usukidoll):

so we already factored.. now we are solving for x. Just split the problems in 2 x-2 = 0, and x +7 =0 solve for x in each case. you're right it's x = 2, -7

OpenStudy (usukidoll):

*in 2 cases.

OpenStudy (anonymous):

Okay thanks :)